OpenStudy (anonymous):

f(t) = [(t-2)/(2t+1)]^9, then f'(t) = ?

OpenStudy (anonymous):

\[ f(t)=\left(\frac{t-2}{2t+1}\right)^9 \] Just use chain rule + quotient rule.

OpenStudy (anonymous):

i used the chain and quotient rule but i don't get one of the MCQ answers

OpenStudy (anonymous):

If you have forgotten:\[ \text{Chain rule:} (f\circ g)'(x)=f'(g(x))\cdot g'(x)\\ \text{Quotient rule:} \left(\frac{f}{g}\right)'(x)=\frac{f'(x)g(x)-f(x)g'(x)}{g^2(x)} \] If you have used those rules and gotten an incorrect answer, please show us your work and result and we can let you know where you made an error.

OpenStudy (anonymous):

here is my working out. @nbouscal

OpenStudy (anonymous):

@remnant i think it's correct .

OpenStudy (anonymous):

That looks correct, what are the MCQ answers? Usually they are just written in a slightly different way and you just have to do a bit of algebraic manipulation to get to one of them.

OpenStudy (anonymous):

I've attached the answers. I've realisd that the answer is A. I just needed to manipulate the answer a little. Thanks for the help. Much appreciated.

OpenStudy (anonymous):

so a is correct .